Show that the vectors (3, 0, -3), (-1, 1, 2), (2, 1, 1) and (4, 2, -2) are linearly dependent in R3.


1
Expert's answer
2020-11-09T20:09:34-0500

Taking into account that "(3, 0, -3)+ (-1, 1, 2)=(2,1,-1)" and "(4,2,-2)=2(2,1,-1)=2[(3, 0, -3)+ (-1, 1, 2)]=2(3, 0, -3)+ 2(-1, 1, 2)=2(3, 0, -3)+ 2(-1, 1, 2)+0(2,1,1)"

we conclude that the vector "(4,2,-2)=2(3, 0, -3)+ 2(-1, 1, 2)+0(2,1,1)" is a linear combimations of vectors "(3, 0, -3), (-1, 1, 2)" and "(2,1,1)", and therefore  the vectors "(3, 0, -3), (-1, 1, 2), (2, 1, 1)" and "(4, 2, -2)" are linearly dependent in "R^3" .